# Conformation and dynamics of wet tangentially-driven active filaments

### v1

@@ III Conformational features of wet active polymers
We first visually inspect the effects of HI on the conformation of the chains. Fig. 2 shows typical representative snapshots of active polymers with and without HI for two bending stiffness values $\kappa=0$ and 10 at low and high active forces given by $f^{a}=0.1$ and 25. With the increase of activity level, both wet and dry active polymers exhibit more shrunken conformations at the higher active force of $f^{a}=25$, with the wet active polymers appearing slightly more shrunken. Moving to the semiflexible chains with $\kappa=10$, the overall conformation of the dry and the wet chains at the low activity level of $f^{a}=0.1$ look alike. In contrast, at high active force of $f^{a}=25$, fluid-mediated interactions induce a notable shrinkage in wet active polymers compared to their dry counterparts.
@@ III.1 End-to-end distance
To have a quantitative assessment of hydrodynamic effects on the mean size of active polymers, we examine the root of the mean squared end-to-end distances of the chains, i.e., $\sqrt{\langle R_{e}^{2}\rangle}=\sqrt{\langle|\mathbf{\bm{r}}_{N}-\mathbf{\bm{r}}_{1}|^{2}\rangle}$. Fig. 3 depicts the $\sqrt{\langle R_{e}^{2}\rangle}$ against activity for dry and wet chains with $\kappa=0$ and $10$. For flexible chains with $\kappa=0$, we observe slightly smaller mean end-to-end distance values for wet polymers compared to their dry counterparts. As the activity level increases, the difference between the two becomes more pronounced. Looking into the $\sqrt{\langle R_{e}^{2}\rangle}$ of semiflexible chains with $\kappa=10$ in Fig. 3, we observe a more notable shrinkage of chains compared to their flexible counterparts, specially at larger active forces $f^{a}>1$ where the mean size of wet active polymers is about twice smaller than the dry ones.
@@ III.2 Interachain correlations
The bond-bond correlation function is defined as $\langle cos(\theta(s))\rangle$, where $\theta(s)$ is the angle between two bond vectors with a curvilinear distance $s$. Fig. 6 shows the bond-bond correlation functions of wet and dry chains with low and high activities at two different degrees of flexibility. For flexible polymers with $\kappa=0$, the correlations drop to zero at small curvilinear distances ($s<5$), regardless of the activity of the chains or inclusion of HI, see Fig. 6(a) and (b). However, at the high active force of $f^{a}=25$, $\langle cos(\theta(s))\rangle$ of wet active chains exhibits a slightly negative dip reflecting anti-correlation of bond vectors at curvilinear distances around $s\approx 10$ and chain folding. For semiflexible chains with $\kappa=10$, at the lower activity level with $f^{a}=0.1$, the dry and wet polymers exhibit very similar bond-bond correlation functions with the $\langle cos(\theta(s))\rangle$ decaying in slightly shorter distances than their dry counterparts. The impact of HI is most remarkable at the high activity of $f^{a}=25$, shown in Fig. 6(d), where the bond-bond correlation of the wet and dry active chains is notably different. While the bond correlation of the dry active chain decays monotonically to zero, the bond-bond correlation functions of wet active polymers exhibits a negative cusp at short curvilinear distances ($s\approx 10$ comparable to their bare persistence length $\ell_{p}^{0}=10$) reflecting the folding of active polymers at this scale resulting in helix-like coiling of chains as can be seen from Fig. 2 (g).
@@
To quantify the combined effects of activity and HI on effective stiffness of the polymer backbone, we obtain an effective persistence length $\ell_{p}$ defined with the following protocol. First, we define $s_{e}$ as the shortest curvilinear distance at which the correlation becomes smaller than $e^{-1}$, i.e, $\langle cos(\theta(s_{e}))\rangle\leq e^{-1}$. We then fit the bond correlation data in the range $0\leq s\leq s_{e}$ by an exponential function $\langle cos(\theta(s))\rangle=\exp(-s/\ell_{p})$ as shown in Fig. 6 and evaluate the persistence length. Fig. 7 shows the effective persistence length against activity for dry and wet chains with $\kappa=0$ and $10$.
We recall that according to the worm-like chain model Binder et al. (2020), the bare persistence length of passive dry chains is $\ell_{p}^{0}=\kappa\sigma/(k_{B}T)$.
@@
As can be seen from Fig. 7, for flexible chains with $\kappa=0$, effective persistence lengths of wet and dry active chains are identical and almost independent of their activity levels. For semiflexible chains with $\kappa=10$, the effective persistence length $\ell_{p}$ of the wet chains are shorter than their dry counterparts. For dry active polymers, $\ell_{p}$ at the lowest activity ($f^{a}=0.01$) is close to the bare persistence length $\ell_{p}^{0}=10$ and that of the corresponding passive chains, see the red dashed line in Fig. 7. The effective persistence length of wet active chains at $f^{a}=0.01$ also agrees with the $\ell_{p}$ of passive wet polymers, as shown by the blue dashed line in Fig. 7. However, its value for the wet passive polymers is slightly smaller than $\ell_{p}^{0}$, reflecting the minor but visible effects of hydrodynamic interactions on the persistence length in the passive limit. The effective persistence lengths decrease with activity for both dry and wet chains, because high active forces increase the likelihood of the polymer backbone overcoming bending energy barriers to adopt more flexible conformations. Consistent with the results of the end-to-end distance, the $\ell_{p}$ of wet semiflexible chains exhibits a significant decrease at high activity levels. For instance, at a high active force of $f^{a}=25$, the effective persistence length of wet semiflexible chains matches that of flexible wet active polymers with $\kappa=0$. This finding, along with the end-to-end distance results, underscore the significant impact of HI on the conformation of highly active semiflexible chains.
@@ III.3 Investigation of mechanism of shrinkage
As we have already seen from Figs. 2(g) and 3, HI lead to marked shrinkage of semiflexible polymers at high activity levels. To understand the underlying mechanism of shrinkage, we investigate the flow field in the vicinity the backbone of an active chain with $\kappa=10$ and $f^{a}=25$ starting from initial conformation of straight line.
Fig. 8 shows snapshots of this chain and the flow field at the monomer positions and their vicinity in a time series. The flow field is measured as the average velocity of MPC particles within a sphere of radius $3$ centered at the position of interest. At $t=0$ when the chain is in a rod-like conformation, we observe a flow field nearly parallel to the backbone but heterogeneous in strength due to intrinsic fluid fluctuations. This heterogeneity in flow field strength, combined with entropic effects, renders the straight conformation unstable. In this example, the flow is stronger near the tail compared to the front segments. Because the back segments move faster than the front segments, they create a compressional flow which pushes the rear segemnts towards the front ones resulting in chain buckling, as illustrated in Fig.8(c) and (d). Finally, the difference in fluid velocity between the buckled and straight regions of the polymer further compresses its conformation until the entire structure becomes helix-like, as seen in Fig. 8(e) and (f). The general pattern in the motion of semiflexible wet active polymers consists of winding helix-like structures induced by flow heterogeneity, which creates local compressional flows and buckles the active polymer locally and create helix-like conformations. This is followed by unwinding of helix due fluctuations which make it unstable due to the resulting extensional flows, see the video in SM for another example.
@@ IV.1 Orientational dynamics
We start by examining the effects of HI on orientational dynamics of active polymers. Since a tangentially-driven polymer is polar (head-tail asymmetry), we define the end-to-end vector as $\mathbf{\bm{R}}_{e}(t)=\mathbf{\bm{r}}_{N}(t)-\mathbf{\bm{r}}_{1}(t)$ where $\mathbf{\bm{r}}_{N}$ corresponds to the head of polar active polymer. In tangentially-driven active polymers, the net active self-propulsion force on the center of mass is proportional to the end-to-end vector, $\mathbf{\bm{F}}^{a}(t)=\sum_{i=1}^{N}\mathbf{\bm{f}}_{i}^{a}(t)=f^{a}\vec{R}_{e}(t)/\ell$. Thus, we characterize the orientational dynamics of the polymers by the normalized time auto-correlation function (TACF) of their end-to-end vector,
@@
To quantify the influence of HI on the decay times of orientational dynamics, we introduce a characteristic orientational relaxation time, $\tau_{e}$, by the following procedure. Firstly, we define $t_{e}$ as the shortest lag time at which the normalized TACF becomes equal to $e^{-1}$. Subsequently, we fit the data within the range $0\leq t\leq t_{e}$ using an exponential function $C_{e}(t)=\text{exp}(-t/\tau_{e})$, shown by dashed lines in Fig. 9, to determine the orientational relaxation time, $\tau_{e}$. Figure 10 presents the extracted orientational relaxation times versus active force for both wet and dry polymers with flexible and semiflexible backbones. Consistent with our previous observations, we find no significant difference in $\tau_{e}$ among chains with varying flexibility levels. However, wet active polymers exhibit notably shorter relaxation times, nearly one order of magnitude smaller, indicating the rapid decay of their end-to-end vector time autocorrelation function (TACF). Additionally, the relaxation times of wet chains follow a $1/f^{a}$ scaling behavior, akin to the findings of previous studies of dry tangentially-driven active polymers Bianco et al. (2018); Fazelzadeh et al. (2023).
@@ IV.2 Translational dynamics
Next we explore the effects of HI on the translational dynamics of active polymers by computing the mean squared displacement (MSD) of the center of mass. Defining the position of the center of mass at any time as $\mathbf{\bm{R}}_{cm}(t)=\frac{1}{N}\sum\mathbf{\bm{r}}_{i}(t)$, the MSD is calculated as $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle=\langle|\mathbf{\bm{R}}_{cm}(t)-\mathbf{\bm{R}}_{cm}(0)|^{2}\rangle$. Fig. 11 presents MSD curves as functions of the lag time for fully flexible and semiflexible polymers, at two active forces $f^{a}=0.1$ and $25$ with and without HI. In all of the MSD plots, we observe three distinct scaling regimes. At very short lag times ($t<10^{3}$ for $f^{a}=0.1$ and $t<1$ for $f^{a}=25$), the fluctuations of random forces predominate the motion of the center of mass. Governed by thermal diffusion, the MSD curves obey the form $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle=6D_{\text{Passive}}t$. For dry active polymers, $D_{\text{Passive}}(\kappa=0)\approx 2.4\times 10^{-5}$ and $D_{\text{Passive}}(\kappa=10)\approx 2.7\times 10^{-5}$, which agree well with the theoretical prediction for the diffusion coefficient of a passive polymer of chain length $N=50$ without HI, i.e., $D_{\text{Passive}}=k_{B}T/(N\gamma)\approx 2.6\times 10^{-5}$. For the wet active polymers the short-time passive diffusion $D_{\text{Passive}}\approx 2\times 10^{-4}$ is one order of magnitude larger than their dry counterparts, reconfirming importance of HI on the dynamics even in the passive limit.
At intermediate timescales, the influence of activity on MSD becomes visible. At this regime the MSD graphs are super-diffusive, i.e., $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle\sim t^{\alpha}$ with $\alpha>1$. At high active force of $f^{a}=25$, the intermediate regime is ballistic ($\alpha=2$), which means the motion of the center of mass of chains can be roughly interpreted as moving straight with a constant speed in the direction of the sum of the total active force and backflow. As discussed earlier, the total active force $\mathbf{\bm{F}}^{a}(t)=f^{a}\mathbf{\bm{R}}_{e}(t)/\ell$ and net backflow force $\mathbf{\bm{F}}^{b}_{cm}(t)=\sum_{i=1}^{N}\mathbf{\bm{F}}^{b}_{mono}(t)=-\frac{NM}{mN+MN_{m}}\mathbf{\bm{F}}^{a}$ are both proportional to the end-to-end vector. Therefore, we expect that the timescale at which the center of mass moves ballistically is set by the relaxation time of the end-to-end vector, i.e., $\tau_{e}$. At low active force of $f^{a}=0.1$, as we have seen in Fig. 10, the orientational relaxation time is of the same order of the passive relaxation time. Consequently, the thermal fluctuations have enough time to perturb the active directed motion of the polymer resulting in a superdiffusive MSD in the intermediate regime with $21<\alpha<$ 2.
@@ V Conclusions
In this paper, we presented a comparative study of the conformational and dynamical features of single tangentially-driven polymer chains in dilute solutions with and without hydrodynamic interactions (HI). We performed computer simulations using coarse-grained bead-spring models, incorporating fluid-mediated interactions using the MPC method revealing significant role of HI on conformational and dynamical properties of tangentially-driven active polymers.

### v2

@@ III Conformational features of wet active polymers
We first visually inspect the effects of HI on the conformation of the chains. Fig. 2 shows typical representative snapshots of active polymers with and without HI for two bending stiffness values $\kappa=0$ and 10 at low and high active forces given by $f^{a}=0.1$ and 25 (see videos S1-S8). With the increase of activity level, both wet and dry active polymers exhibit more shrunken conformations at the higher active force of $f^{a}=25$, with the wet active polymers appearing slightly more shrunken. Moving to the semiflexible chains with $\kappa=10$, the overall conformation of the dry and the wet chains at the low activity level of $f^{a}=0.1$ look alike. In contrast, at high active force of $f^{a}=25$, fluid-mediated interactions induce a notable shrinkage in wet active polymers compared to their dry counterparts.
@@ III.1 End-to-end distance
To have a quantitative assessment of hydrodynamic effects on the mean size of active polymers, we examine the root of the mean squared end-to-end distances of the chains, i.e., $\sqrt{\langle R_{e}^{2}\rangle}=\sqrt{\langle|\mathbf{\bm{R}}_{N_{m}}-\mathbf{\bm{R}}_{1}|^{2}\rangle}$. Fig. 3 depicts the $\sqrt{\langle R_{e}^{2}\rangle}$ against activity for dry and wet chains with $\kappa=0$ and $10$. For flexible chains with $\kappa=0$, we observe slightly smaller mean end-to-end distance values for wet polymers compared to their dry counterparts. As the activity level increases, the difference between the two becomes more pronounced. Looking into the $\sqrt{\langle R_{e}^{2}\rangle}$ of semiflexible chains with $\kappa=10$ in Fig. 3, we observe a more notable shrinkage of chains compared to their flexible counterparts, specially at larger active forces $f^{a}>1$ where the mean size of wet active polymers is about twice smaller than the dry ones.
@@ III.2 Interachain correlations
The bond-bond correlation function is defined as $\langle cos(\theta(s))\rangle$, where $\theta(s)$ is the angle between two bond vectors with a curvilinear distance $s$. Fig. 6 shows the bond-bond correlation functions of wet and dry chains with low and high activities at two different degrees of flexibility. For flexible polymers with $\kappa=0$, the correlations drop to zero at small curvilinear distances ($s<5$), regardless of the activity of the chains or inclusion of HI, see Fig. 6(a) and (b). However, at the high active force of $f^{a}=25$, $\langle cos(\theta(s))\rangle$ of wet active chains exhibits a slightly negative dip reflecting anti-correlation of bond vectors at curvilinear distances around $s\approx 10$ and chain folding. For semiflexible chains with $\kappa=10$, at the lower activity level with $f^{a}=0.1$, the dry and wet polymers exhibit very similar bond-bond correlation functions with the $\langle cos(\theta(s))\rangle$ decaying in slightly shorter distances than their dry counterparts. The impact of HI is most remarkable at the high activity of $f^{a}=25$, shown in Fig. 6(d), where the bond-bond correlation of the wet and dry active chains are notably different. While the bond correlation of the dry active chain decays monotonically to zero, the bond-bond correlation functions of wet active polymers exhibits a negative cusp at short curvilinear distances ($s\approx 10$ comparable to their bare persistence length $\ell_{p}^{0}=10$) reflecting the folding of active polymers at this scale resulting in helix-like coiling of chains as can be seen from Fig. 2 (g).
@@
To quantify the combined effects of activity and HI on effective stiffness of the polymer backbone, we obtain an effective persistence length $\ell_{p}$ defined with the following protocol. First, we define $s_{e}$ as the shortest curvilinear distance at which the correlation becomes smaller than $e^{-1}$, i.e, $\langle cos(\theta(s_{e}))\rangle\leq e^{-1}$. We then fit the bond correlation data in the range $0\leq s\leq s_{e}$ by an exponential function $\langle cos(\theta(s))\rangle=\exp(-s/\ell_{p})$ as shown in Fig. 6 and evaluate the persistence length. Fig. 7 shows the effective persistence length against activity for dry and wet chains with $\kappa=0$ and $10$. We recall that according to the worm-like chain model Binder et al. (2020), the bare persistence length of passive dry chains is $\ell_{p}^{0}=\kappa\sigma/(k_{B}T)$.
@@
As can be seen from Fig. 7, for flexible chains with $\kappa=0$, the effective persistence lengths of wet and dry active chains are identical and almost independent of their activity levels. For semiflexible chains with $\kappa=10$, the effective persistence length $\ell_{p}$ of the wet chains are shorter than their dry counterparts. For dry active polymers, $\ell_{p}$ at the lowest activity ($f^{a}=0.01$) is close to the bare persistence length $\ell_{p}^{0}=10$ and that of the corresponding passive chains, see the red dashed line in Fig. 7. The effective persistence length of wet active chains at $f^{a}=0.01$ also agrees with the $\ell_{p}$ of passive wet polymers, as shown by the blue dashed line in Fig. 7. However, its value for the wet passive polymers is slightly smaller than $\ell_{p}^{0}$, reflecting the minor but visible effects of hydrodynamic interactions on the persistence length in the passive limit. The effective persistence lengths decrease with activity for both dry and wet chains, because high active forces increase the likelihood of the polymer backbone overcoming bending energy barriers to adopt more flexible conformations. Consistent with the results of the end-to-end distance, the $\ell_{p}$ of wet semiflexible chains exhibits a significant decrease at high activity levels. For instance, at a high active force of $f^{a}=25$, the effective persistence length of wet semiflexible chains matches that of flexible wet active polymers with $\kappa=0$. This finding, along with the end-to-end distance results, underscore the significant impact of HI on the conformation of highly active semiflexible chains.
@@ III.3 Investigation of mechanism of shrinkage
As we have already seen from Figs. 2(g) and 3, HI lead to marked shrinkage of semiflexible polymers at high activity levels. To understand the underlying mechanism of shrinkage, we investigate the flow field in the vicinity of the backbone of an active chain with $\kappa=10$ and $f^{a}=25$, initialized as an straight line.
Fig. 8 shows snapshots of this chain and the flow field at the monomer positions and their vicinity in a time series (see video S9). The flow field is measured as the average velocity of MPC particles within a sphere of radius $3$ centered at the position of interest. At $t=0$ when the chain is in a rod-like conformation, we observe a flow field nearly parallel to the backbone but heterogeneous in strength due to intrinsic fluid fluctuations. This heterogeneity in flow field strength, combined with entropic effects, renders the straight conformation unstable. In this example, the flow is stronger near the tail compared to the front segments. Because the back segments move faster than the front segments, they create a compressional flow which pushes the rear segemnts towards the front ones resulting in chain buckling, as illustrated in Fig.8(c) and (d). Finally, the difference in fluid velocity between the buckled and straight regions of the polymer further compresses its conformation until the entire structure becomes helix-like, as seen in Fig. 8(e) and (f). The general pattern in the motion of semiflexible wet active polymers consists of winding helix-like structures induced by flow heterogeneity, which creates local compressional flows and buckles the active polymer locally and create helix-like conformations. This is followed by unwinding of helix due fluctuations which make it unstable due to the resulting extensional flows, see video S8.
@@ IV.1 Orientational dynamics
We start by examining the effects of HI on orientational dynamics of active polymers. Since a tangentially-driven polymer is polar (head-tail asymmetry), we define the end-to-end vector as $\mathbf{\bm{R}}_{e}(t)=\mathbf{\bm{R}}_{N_{m}}(t)-\mathbf{\bm{R}}_{1}(t)$ where $\mathbf{\bm{R}}_{N_{m}}$ corresponds to the head of polar active polymer. In tangentially-driven active polymers, the net active self-propulsion force on the center of mass is proportional to the end-to-end vector, $\mathbf{\bm{F}}^{a}(t)=\sum_{i=1}^{N}\mathbf{\bm{f}}_{i}^{a}(t)=f^{a}\vec{R}_{e}(t)/\ell$. Thus, we characterize the orientational dynamics of the polymers by the normalized time auto-correlation function (TACF) of their end-to-end vector,
@@
To quantify the influence of HI on the decay times of orientational dynamics, we introduce a characteristic orientational relaxation time, $\tau_{e}$, by the following procedure. First, we define $t_{e}$ as the shortest lag time at which the normalized TACF becomes equal to $e^{-1}$. Subsequently, we fit the data within the range $0\leq t\leq t_{e}$ using an exponential function $C_{e}(t)=\text{exp}(-t/\tau_{e})$, shown by dashed lines in Fig. 9, to determine the orientational relaxation time, $\tau_{e}$. Figure 10 presents the extracted orientational relaxation times versus active force for both wet and dry polymers with flexible and semiflexible backbones. Consistent with our previous observations, we find no significant difference in $\tau_{e}$ among chains with varying flexibility levels. However, wet active polymers exhibit notably shorter relaxation times, nearly one order of magnitude smaller, indicating the rapid decay of their end-to-end vector time autocorrelation function (TACF). Additionally, the relaxation times of wet chains follow a $1/f^{a}$ scaling behavior, akin to the findings of previous studies of dry tangentially-driven active polymers Bianco et al. (2018); Fazelzadeh et al. (2023).
@@ IV.2 Translational dynamics
Next we explore the effects of HI on the translational dynamics of active polymers by computing the mean squared displacement (MSD) of the center of mass. Defining the position of the center of mass at any time as $\mathbf{\bm{R}}_{cm}(t)=\frac{1}{N_{m}}\sum\mathbf{\bm{R}}_{i}(t)$, the MSD is calculated as $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle=\langle|\mathbf{\bm{R}}_{cm}(t)-\mathbf{\bm{R}}_{cm}(0)|^{2}\rangle$. Fig. 11 presents MSD curves as functions of the lag time for fully flexible and semiflexible polymers, at two active forces $f^{a}=0.1$ and $25$ with and without HI. In all of the MSD plots, we observe three distinct scaling regimes. At very short lag times ($t<10^{3}$ for $f^{a}=0.1$ and $t<1$ for $f^{a}=25$), the fluctuations of random forces predominate the motion of the center of mass. Governed by thermal diffusion, the MSD curves obey the form $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle=6D_{\text{Passive}}t$. For dry active polymers, $D_{\text{Passive}}(\kappa=0)\approx 2.4\times 10^{-5}$ and $D_{\text{Passive}}(\kappa=10)\approx 2.7\times 10^{-5}$, which agree well with the theoretical prediction for the diffusion coefficient of a passive polymer of chain length $N=50$ without HI, i.e., $D_{\text{Passive}}=k_{B}T/(N\gamma)\approx 2.6\times 10^{-5}$. For the wet active polymers the short-time passive diffusion $D_{\text{Passive}}\approx 2\times 10^{-4}$ is one order of magnitude larger than their dry counterparts, reconfirming importance of HI on the dynamics even in the passive limit.
At intermediate timescales, the influence of activity on MSD becomes visible. At this regime the MSD graphs are super-diffusive, i.e., $\langle\Delta\mathbf{\bm{R}}^{2}_{cm}(t)\rangle\sim t^{\alpha}$ with $\alpha>1$. At high active force of $f^{a}=25$, the intermediate regime is ballistic ($\alpha=2$), which means the motion of the center of mass of the chains can be roughly interpreted as moving straight with a constant speed in the direction of the sum of the total active force and backflow. As discussed earlier, the total active force $\mathbf{\bm{F}}^{a}(t)=f^{a}\mathbf{\bm{R}}_{e}(t)/\ell$ and net backflow force $\mathbf{\bm{F}}^{b}_{cm}(t)=\sum_{i=1}^{N}\mathbf{\bm{F}}^{b}_{mono}(t)=-\frac{NM}{mN+MN_{m}}\mathbf{\bm{F}}^{a}$ are both proportional to the end-to-end vector. Therefore, we expect that the timescale at which the center of mass moves ballistically is set by the relaxation time of the end-to-end vector, i.e., $\tau_{e}$. At low active force of $f^{a}=0.1$, as we see in Fig. 10, the orientational relaxation time is of the same order of the passive relaxation time. Consequently, the thermal fluctuations have enough time to perturb the active directed motion of the polymer resulting in a superdiffusive MSD in the intermediate regime with $21<\alpha<$ 2.
@@ V Conclusions
In this paper, we presented a comparative study of the conformational and dynamical features of single tangentially-driven polymer chains in dilute solutions with and without hydrodynamic interactions (HI). We performed computer simulations using coarse-grained bead-spring models, incorporating fluid-mediated interactions using the MPCD method, revealing significant role of HI on conformational and dynamical properties of tangentially-driven active polymers.